OpenStudy (anonymous):

Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240lb mixture that is 8% protein? How many pounds of cornmeal? How many pounds of soybean?

OpenStudy (anonymous):

assume x is the amount of soyabean , whilst y is the amount of cornmeal taken. so u hav 2 equations, x+y=240 and 12x+6y=8; linear equations , solve them to get the answer:

OpenStudy (anonymous):

get it ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i need sleep but im also trying to do more than I can rihgtn ow I guess, I did this same problem the other day lol

OpenStudy (anonymous):

oh, chill now, rest for sometime and then read it. see the total weight of the food has to be 240lb rite? so, the total weight of the composite food x+y,has to be 240. now the protein content of the food is fixed @ 8%. therefore wen u include x amount of soyabean, the protein contribution from it is, 12% of x. similar is for cornmeal. now this individual contribution has to be equal to 8% . hence the equations formed :)

OpenStudy (anonymous):

8% of 240 = 19.2lb Now, let amount of soybean is x and that of cornmeal is 240-x 12% of x = x*12/100 = 3x/25 And, 6% of (240-x) = (240-x)*6/100 = (240-x)*3/50 So, 3x/25 + (240-x)*6/100 = 19.2 Solving this equation, x = 80 So, amount of soybean is 80 lb and that of cornmeal is 160lb

OpenStudy (anonymous):

the second equation i've given has to be 12x+6y=8*240 : sorry :)